if two vectors P and Q acting at a point have the resultant of magnitude root 3Q making an angle 30 degree with the direction of P then P:Q is?

Grade:11

1 Answers

dhruba
2 years ago
i hope u could make the figure there u can see that the angle between p vector and resultant vector equals 30 degree as given now complete the parallelogram and u get a traingle let the angle between r vector and q vector =theta  ….then angle between p vector and q vector =150-theta )angle sum property) ….now just apply the sine rule so 
you get sin 30^/q=sin theta/p =sin(150-theta)/root3q
 
then u get 
p/q=2sin theta
 
and sin(150-theta)=root3/2
so.150-theta=60 ….theta=90^
 
therefore p/q=2sin theta=2 x (sin 90^)=2